# GS-ProCams: Gaussian Splatting-based Projector-Camera Systems

### v1

@@ 3.4 Lighting
Projector direct light. Denote the projector gamma as $\gamma_{p}$, which converts standard sRGB color into linear RGB space. Denote the projector’s luminous power as $G_{p}$, and assume constant light attenuation. Additionally, due to the projector’s optical characteristics, e.g., point spread function (PSF) [38], the projected light suffers from defocus. Therefore, we use a learnable $5\times 5$ kernel $\kappa$ to model the projector’s PSF. Consequently, the projector emitted direct light $L_{p}$ can be expressed as:
@@
where $\mathbf{x}_{p}$ is the projector pixel that directly illuminates the surface point $\mathbf{x}_{s}$, as shown in Fig. 2, and $\ast$ is the convolution operator. Denote the projector projection matrix as $\mathbf{M}_{p}$, the relationship between $\mathbf{x}_{p}$ and $\mathbf{x}_{s}$ is given by
@@
where $\overline{\mathbf{x}}_{p}$ and $\overline{\mathbf{x}}_{s}$ denote the homogeneous coordinates of $\mathbf{x}_{p}$ and $\mathbf{x}_{s}$, respectively. Then, according to Eq. 4, the outgoing/reflected direct light at $\mathbf{x}_{s}$ can be computed by:
@@
where the normal $\mathbf{n}_{s}$ of the surface point $\mathbf{x}_{s}$ is determined by Eq. 6. We approximate the BRDF $f$ in Eq. 10 using a simplified Disney BRDF model [8].
Global illumination. In most ProCams settings, projector direct light dominates the illumination, therefore, we utilize a simplified generalized estimate to replace the precise yet complex calculations for efficiency. Inspired by 3DGS [35] and 2DGS [27], we approximate the global illumination radiance as a view-dependent color $\mathbf{c}_{r}(\omega_{o})$ by maintaining a set of coefficients of the spherical harmonics for each Gaussian point and then rasterize the color using alpha blending according to Eq. 3 as follows:
@@
Finally, the camera-captured color is given by
$\hat{\mathbf{I}}_{c}=\text{clamp}\left((\mathbf{C}_{p}+\mathbf{C}_{g})^{\frac{1}{\gamma_{c}}},0,1\right),$ (12)
@@
Figure 4: Visual comparisons of ProCams simulation on the Nepmap synthetic dataset [14]. The first column displays an object from a novel viewpoint, the second column shows the object under a novel projection pattern, and the third and fourth columns present the results of two methods. We present two of the four scenes here as examples. Compared to Nepmap, our model exhibits finer details and more realistic color. Moreover, our method outperforms Nepmap on computation and memory efficiency by a significant margin (Table 2). See supplementary material for more results.
@@ 3.5 Optimization
As shown in Fig. 3, the optimizable parameters associated with each Gaussian point are $\{\mathbf{p},\mathbf{q},\mathbf{s},o,\mathbf{c}_{g},\mathbf{b},r\}$. The optimizable parameters of projector responses are $\{\gamma_{p},G_{p},\kappa\}$. We impose the following constraints to optimize the GS-ProCams parameters.
Photometric loss. In line with 3DGS [35], we consider the combined loss function of $\mathcal{L}_{1}$ and D-SSIM $\mathcal{L}_{D-SSIM}$ to minimize the photometric difference between the simulated result $\hat{\mathbf{I}}_{c}$ and camera captured ground truth $\mathbf{I}_{c}$ and set the weight $\lambda=0.2$:
$\mathcal{L}_{c}=(1-\lambda)\mathcal{L}_{1}+\lambda\mathcal{L}_{D-SSIM}.$ (13)
@@
where $\mathbf{n}_{i}$ is the inherent normal of the the intersection between $\mathbf{x}$ and the $i$-th splat.
Materials regularization. In this paper, we estimate surface materials from realistic scenes where roughness maps are anticipated to exhibit smooth characteristics in regions with uniform color [79]. Consequently, we define a bilateral smoothness term, guided by the albedo, as follows:
$\mathcal{L}_{material}={\left\|\nabla{R}\right\|}\text{exp}(-{\left\|\nabla{\mathbf{B}}\right\|}).$ (15)
Note that we detach the gradient propagation of the albedo $\mathbf{B}$ before computing this loss.
Mask entropy. Most projector applications focus only on regions of interest, such as the projector’s FOV [26] or objects [14], typically using binary masks to indicate these areas. In light of this, We introduce a cross-entropy constraint [71] to regularize the optimization for obtaining opaque surfaces by forcing the accumulated opacity $O=\sum_{i\in K}W_{i}$ to be aligned with the mask $O_{m}$. The constraint is defined as:
$\mathcal{L}_{entropy}=-O_{m}\log O-(1-O_{m})\log(1-O),$ (16)
@@
$\mathcal{L}=\mathcal{L}_{c}+\lambda_{d}\mathcal{L}_{d}+\lambda_{n}\mathcal{L}_{n}+\lambda_{m}\mathcal{L}_{material}+\lambda_{e}\mathcal{L}_{entropy},$ (17)
@@
where $\lambda_{d}$, $\lambda_{n}$, $\lambda_{m}$, and $\lambda_{e}$ represent the weights for the respective loss terms, which are set to $1000$, $0.05$, $0.002$ and $0.1$, respectively. We jointly optimize the GS-ProCams parameters using Adam optimizer [36] for 30,000 training steps.
@@ 4 Experiments
To demonstrate the advantages of our approach, we compared it with several state-of-the-art ProCams simulation and/or projector compensation methods [14, 24, 26] in Sec. 4.1 and Sec. 4.2. In our experiments, we employed Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) [82] to evaluate the ProCams simulation, and projector compensation quality. Additionally, we statistically examined the memory requirements, training time, and inference frames per second (FPS) on an NVIDIA A800 GPU.
@@ 4.1 Synthetic dataset
We compared our GS-ProCams with the state-of-the-art NeRF-based ProCams simulation approach [14], using their synthetic dataset to evaluate the ProCams simulation quality. We refer to this method as Nepmap for convenience. We utilized the four synthetic scenes released by Nepmap and their pre-trained models to compare 36 novel viewpoints and novel projection patterns for each scene. Quantitative results are presented in Tab. 2, and qualitative comparisons are illustrated in Fig. 4. The results indicate that our model not only outperforms Nepmap in terms of quality but also shows a significant improvement in efficiency, i.e., about 600 times faster and only uses 1/10 of GPU memory. Furthermore, our model does not require additional devices and takes ambient light into account as listed in Tab. 1.
Table 2: ProCams simulation results on the Nepmap synthetic dataset [14]. Results are averaged over 4 different scenes.
@@
Nepmap [14] 27.23 0.945 0.061 $>3\text{hrs}$ 0.36 34.88 GB
GS-ProCams 33.07 0.974 0.032 9.3 min 222 02.02 GB
Figure 5: Visual comparisons of ProCams simulation on real-world benchmark dataset. We compare with DeProCams [24] (view-specific) at the last training viewpoint (first row) and at the novel viewpoint that our GS-ProCams has not previously seen (second row). Note that DeProCams (view-specific) necessitates repeated data acquisition and training for the novel viewpoint, whereas our approach can generalize to novel viewpoints. The first column displays the projector input pattern, the second column shows the surface (no superimposed projection) from the viewpoint, and the subsequent columns sequentially illustrate the camera-captured ground truth and the simulations results. See supplementary material for more results.
Figure 6: Intermediate results. GS-ProCams uses interpretable parameters to establish the geometric and photometric mappings of ProCams, jointly estimating them while achieving view-agnostic ProCams simulation. Here, “$L_{p}$” is in the linear color space. “Albedo” and “Global illumination” are converted from linear to sRGB color space for better visual experience.
Table 3: ProCams simulation comparisons on real-world benchmark dataset. Results are averaged over 10 different scenes. Note that DeProCams needs to additionally capture training data and retraining for novel viewpoints, while our GS-ProCams waives this requirement.
Method Trained Viewpoint Unseen Novel Viewpoint
[…]
DeProCams11footnotemark: 1 [24] 28.66 0.907 0.109 04.61 min 27.63 0.886 0.134 4.62 min 26.73 GB
w/o PSF (Ours)22footnotemark: 2 25.76 0.887 0.104 15.86 min 24.83 0.837 0.121 0 03.38 GB
GS-ProCams (Ours)22footnotemark: 2 27.22 0.911 0.083 15.69 min 26.11 0.859 0.100 0 03.49 GB
11footnotemark: 1 View-specific method; 22footnotemark: 2 View-agnostic method
4.2 Real-world benchmark dataset
We collected a benchmark dataset in real indoor lighting environments. Our projector-camera system consists of a Panasonic Lumix DC-S5 camera and an EPSON CB-965 projector, and an RGB fill light was employed to create various ambient lighting conditions. It is worth noting that unlike Nepmap [14], our method does not require additional light sources to work, the RGB fill light is only used to change ambient lighting conditions. The projector and camera resolutions are both set to $800\times 800$.
We projected multiple patterns for each viewpoint, including a fixed pattern to enhance surface texture for the SfM process [58] and a natural image for model training. We followed the methodology outlined in [26] to obtain projector FOV masks. In our experimental configuration, we projected natural patterns across 48 viewpoints. At the 48th viewpoint, an additional 48 natural patterns were projected for validation. Subsequently, we randomly repositioned the camera and projected 48 patterns once more to evaluate the applicability of our method from this novel viewpoint. See supplementary material for more details.
Because Nepmap [14] requires a light source to be co-located with the camera and only works for a dark room, we cannot compare with it on our ProCams settings. Instead, we compare with two state-of-the-art view-specific methods [24, 26].
Figure 7: Visual comparisons of projector compensation on real-world benchmark dataset. Similar to ProCams simulation (Fig. 5), the camera-captured results from different viewpoints within the scene are presented in two rows. Our model is applied directly to novel viewpoints without retraining in the second row, while the other two view-specific methods necessitate repeated data capturing and training. The first three columns respectively illustrate the surface appearance from the specified viewpoint, the uncompensated appearance, and the expected appearance. The last three columns display the compensated results from different models. Note that our method effectively compensates for geometry and photometry when applied to the novel viewpoint. See supplementary material for more results.
Table 4: Projector compensation results on real-world dataset. Results are averaged over 10 different scenes. Our approach is directly applied to the unseen novel viewpoint without retraining, while the other two methods must be retrained.
Method Trained Viewpoint Unseen Novel Viewpoint
[…]
DeProCams11footnotemark: 1 [24] 22.78 0.816 0.189 04.61 min 22.35 0.796 0.217 04.62 min 26.73 GB
CompenNeSt++11footnotemark: 1 [26] 24.02 0.834 0.161 13.25 min 24.05 0.830 0.173 13.21 min 30.79 GB
GS-ProCams (Ours)22footnotemark: 2 23.91 0.828 0.173 15.69 min 23.66 0.807 0.194 0 03.49 GB
Uncompensated 14.80 0.669 0.281 - 14.57 0.641 0.303 - -
11footnotemark: 1 View-specific method; 22footnotemark: 2 View-agnostic method
Figure 8: Ablation study of PSF estimation. We compare GS-ProCams with a degraded version that ignores projector PSF. Note that the result of “w/o PSF” shown in the zoomed-in patches exhibits unrealistically sharp projections.
ProCams simulation. We use the state-of-the-art method, DeProCams [24] (view-specific), as a baseline to evaluate the effectiveness of our approach (view-agnostic) at the unseen novel viewpoint. We present quantitative comparisons in Tab. 3 and illustrate qualitative findings with a scene example in Fig. 5. For the trained viewpoint, our GS-ProCams outperforms DeProCams, due to our GS-based formulation. For the unseen novel viewpoint, DeProCams shows better simulation quality, because it is view-specific and is trained on this viewpoint, while our GS-ProCams has never seen this viewpoint. Our method highlights its advantages compared to the view-specific approach. First, our method facilitates the synthesis of surface appearances with novel projections under novel viewpoints. Second, our method demonstrates a significant efficiency advantage in memory usage. Additionally, as previously noted, GS-ProCams effectively solves the view-agnostic projection mapping problem by physically-based rendering, while deriving interpretable quantities from multi-view images as shown in Fig. 6.
Projector compensation. Our proposed method facilitates ProCams simulation through a fully differentiable process, enabling the optimization of projection patterns to align rendered images with the desired effects. Once trained, our method allows for direct projector compensation from novel viewpoints without retraining or fine-tuning. We compare with two state-of-the-art projector compensation methods CompenNeSt++ [26] (view-specific) and DeProCams [24] (view-specific). The quantitative results are summarized in Tab. 4, while the visual comparisons are illustrated in Fig. 7. The results show that our method outperforms DeProCams under the trained viewpoint and is comparable to CompenNeSt++. For the novel viewpoint, there is a gap between our model and CompenNeSt++, but it still outperforms DeProCams. This is because both CompenNeSt++ and DeProCams are retrained for the novel viewpoint, but our GS-ProCams is directly applied to the novel viewpoint without retraining. The view-agnostic projector compensation and low memory requirements demonstrate our significant advantages in practical applications.

### v2

@@ 2.4 Lighting
Projector direct light. Denote the projector gamma as $\gamma_{p}$, which converts standard sRGB color into linear RGB space. Denote the projector’s luminous power as $G_{p}$, and assume constant light attenuation. Additionally, due to the projector’s optical characteristics, e.g., point spread function (PSF), the projected light suffers from defocus. Therefore, we use a learnable $5\times 5$ kernel $\kappa$ to simply approximate the projector’s PSF. Consequently, the projector emitted direct light $L_{p}$ can be expressed as:
@@
where $\mathbf{x}_{p}$ is the projector pixel that directly illuminates the surface point $\mathbf{x}_{s}$, as shown in Fig.˜1, and $\ast$ is the convolution operator. Theoretically, the PSF varies spatially in the image plane [30, 36], but accurately measuring or modeling this variation is challenging and computationally intensive. For efficiency, we use a spatially-invariant approximation, with some sacrifice in accuracy.
Denote the projector projection matrix as $\mathbf{M}_{p}$, the relationship between $\mathbf{x}_{p}$ and $\mathbf{x}_{s}$ is given by
@@
where $\overline{\mathbf{x}}_{p}$ and $\overline{\mathbf{x}}_{s}$ denote the homogeneous coordinates of $\mathbf{x}_{p}$ and $\mathbf{x}_{s}$, respectively. Then, according to Eq.˜4, the outgoing/reflected projector direct light at $\mathbf{x}_{s}$ can be computed by:
@@
where the normal $\mathbf{n}_{s}$ of the surface point $\mathbf{x}_{s}$ is determined by Eq.˜6. We approximate the BRDF $f$ in Eq.˜10 using a simplified Disney BRDF model [7].
Global illumination. In most projection mapping settings, direct projector illumination dominates surface appearance and is explicitly modeled in our framework. Meanwhile, global illumination includes environment lighting and weaker projector-induced indirect effects, which also contribute to the final appearance. Although projector-induced indirect effects are input-dependent, accurately modeling them is computationally expensive [38]. Inspired by GS-based radiance fields [33, 26], we approximate residual global illumination as a view-dependent residual color $\mathbf{c}_{g}(\omega_{o})$ by maintaining spherical harmonics coefficients for each Gaussian point, which are then rasterized by:
@@
To mitigate degeneracy during joint optimization, we incorporate a plain black projection at training viewpoints to constrain the residual term. This residual primarily captures environmental illumination, while projector-induced indirect effects are coarsely approximated and implicitly absorbed into this term and other components such as albedo during joint optimization. This formulation also supports rendering beyond the projection surface, although performance remains limited in textureless regions of the surrounding scene, which is a known challenge for GS-based radiance methods. Finally, the camera-captured color is given by
$\hat{\mathbf{I}}_{c}=\text{clamp}\left((\mathbf{C}_{p}+\mathbf{C}_{g})^{{\gamma_{c}}},0,1\right),$ (12)
@@
Figure 3: Visual comparisons of ProCams simulation on the Nepmap synthetic dataset [13]. The first column displays an object from a novel viewpoint, the second column shows the object under a novel projection pattern, and the third and fourth columns present the results of two methods. We present two of the four scenes here as examples. Compared to Nepmap, our model exhibits finer details and more realistic colors. Moreover, our method outperforms Nepmap in computation and memory efficiency by a significant margin (Table 2). See supplementary material for more results.
Figure 4: Learned albedo on the Nepmap synthetic dataset [13]. We present the albedo maps learned by Nepmap and our GS-ProCams under different novel viewpoints of the same scene, shown in two separate rows. The albedo maps are converted from linear to sRGB color space for a better visual experience. The first column shows the scene under camera co-located light only. The second column depicts the same scene under a uniform white projection. GS-ProCams models the co-located lighting as part of the residual color and includes projector occlusions within the albedo estimation.
@@ 2.5 Optimization
As shown in Fig.˜2, the optimizable parameters associated with each Gaussian point are $\{\mathbf{p},\mathbf{q},\mathbf{s},o,\mathbf{c}_{g},\mathbf{b},r\}$. The optimizable parameters of projector responses are $\{\gamma_{p},G_{p},\kappa\}$. We impose the following constraints to optimize the GS-ProCams parameters.
Photometric loss. In line with 3DGS [33], we consider the combined loss function of $\mathcal{L}_{1}$ and DSSIM $\mathcal{L}_{\text{DSSIM}}$ to minimize the photometric difference between the simulated result $\hat{\mathbf{I}}_{c}$ and camera captured ground truth $\mathbf{I}_{c}$ and set the weight $\lambda=0.2$:
$\mathcal{L}_{c}=(1-\lambda)\mathcal{L}_{1}+\lambda\mathcal{L}_{\text{DSSIM}}$ (13)
@@
where $\mathbf{n}_{i}$ is the inherent normal of the intersection between $\mathbf{x}$ and the $i$-th splat.
Materials regularization. In this paper, we jointly estimate surface materials from realistic scenes where roughness maps are anticipated to exhibit smooth characteristics in regions with uniform color [78]. Consequently, we define a bilateral smoothness term, guided by the albedo, as follows:
$\mathcal{L}_{\text{material}}={\left\|\nabla{R}\right\|}\text{exp}(-{\left\|\nabla{\mathbf{B}}\right\|}).$ (15)
Note that we detach the gradient propagation of albedo $\mathbf{B}$ before computing this loss.
Mask entropy. Most projector applications focus only on regions of interest, such as the projector’s FOV [24] or objects [13], typically using binary masks to indicate these areas. Therefore, we introduce a cross-entropy constraint [71], to regularize the optimization for obtaining opaque projection surfaces when $O_{m}$ is available, and to prune Gaussians outside the region of interest for efficiency. The constraint forces the accumulated opacity $O=\sum_{i\in K}W_{i}$ to align with $O_{m}$ and is defined as:
$\mathcal{L}_{\text{entropy}}=-O_{m}\log O-(1-O_{m})\log(1-O),$ (16)
@@
$\mathcal{L}=\mathcal{L}_{c}+\lambda_{d}\mathcal{L}_{d}+\lambda_{n}\mathcal{L}_{n}+\lambda_{m}\mathcal{L}_{\text{material}}+\lambda_{e}\mathcal{L}_{\text{entropy}},$ (17)
@@
where $\lambda_{d}$, $\lambda_{n}$, $\lambda_{m}$, and $\lambda_{e}$ are the weights for the respective loss terms, selected as 1000, 0.05, 0.002, and 0.1, respectively, based on prior work [26, 16] for robust optimization. We jointly optimize the GS-ProCams parameters using Adam optimizer [34] for 20,000 training steps.
Figure 5: Visual comparisons of ProCams simulation on real-world benchmark dataset. We compare with DeProCams [23] (view-specific) and DPCS [38] (view-specific), for novel projection synthesis. The first row shows a viewpoint included in GS-ProCams training, while the second row shows a novel viewpoint from the same scene that GS-ProCams has never seen during training. The first column shows the same projector input pattern, the second column presents the projection surface under a uniform gray projection and captured in different camera viewpoints, and the remaining columns show the camera-captured ground truth and simulation results. Notably, view-specific methods necessitate repeated data acquisition and training, even for the novel viewpoint, whereas our approach generalizes to novel viewpoints without seeing/retraining on them. Our GS-ProCams consistently produces high-quality simulation results for the novel projection under both trained and novel viewpoints. See supplementary material for more results.
2.6 Projector compensation
Our GS-ProCams facilitates ProCams simulation through a fully differentiable process, enabling the optimization of projector input patterns to align rendered images with the desired appearances. Once trained, GS-ProCams represents projector compensation as an inverse rendering problem. Given the desired appearance in the camera frame as $\mathbf{I}_{d}$, our framework allows for optimizing a virtual projector input pattern $\hat{\mathbf{I}}_{p}$ to make the simulation result $\hat{\mathbf{I}}_{c}$ in the virtual space align with $\mathbf{I}_{d}$ as follows:
where $\hat{\mathbf{I}}^{*}_{p}$ is the compensated projector input for the desired visual projection appearance $\mathbf{I}_{d}$.
Table 2: ProCams simulation results on the Nepmap synthetic dataset [13]. Results are averaged over 4 scenes. Since synthetic scenes ignore defocus blur, we use a degraded version of GS-ProCams that excludes the projector PSF for these experiments.
[…]
Nepmap [13] 27.24 / 0.9445 / 0.061 $\sim$3 hrs >>24 GB 0.3
GS-ProCams w/o PSF 32.12 / 0.9695 / 0.036 3.9 min 2.03 GB 275
Adopted from the original paper
Figure 6: Intermediate results. GS-ProCams explicitly models the ProCams’ geometric and photometric mapping using interpretable parameters. On the left, we show the projection surface (Surface-View) and the scene under environment light only (Environment). On the right, are GS-ProCams learned intermediate parameters, $L_{p}$” is in the linear color space (thus looks overexposed), while “Albedo” and “Residual color” are converted from linear to sRGB for better visual presentation.
Table 3: ProCams simulation comparisons on a real-world benchmark dataset. Results are averaged over 7 different scenes. For each scene, we compare our GS-ProCams with baseline methods using 5 training viewpoints and 8 additional novel viewpoints. At each viewpoint, 25 novel projections are evaluated. Note that view-dependent methods must capture additional training data and retrain the model for each single viewpoint, regardless of training or novel viewpoints, while our method only need to be trained on the training viewpoints and can generalize to unseen novel viewpoints without retraining. “# Train” indicates the number of textured patterns used for training each method. $\mathbf{d}_{\text{err}}$ denotes the mean $L_{2}$ distance between the inferred and the manually cleaned ground truth point cloud obtained by SL [44]. We use image masks obtained from SL for both inferred and ground truth images to ensure consistency.
Trained Viewpoints Novel Viewpoints
[…]
100 DeProCams [23] 25.07 0.9216 0.079 0.1106 25.12 0.9134 0.085 0.1264
DPCS [38] 26.17 0.9354 0.050 - 25.87 0.9290 0.055 -
GS-ProCams w/o PSF 24.25 0.9242 0.064 0.0697 23.74 0.9117 0.072 0.0611
GS-ProCams (Ours) 26.76 0.9444 0.045 0.0670 26.23 0.9341 0.051 0.0556
75 DeProCams [23] 24.55 0.9158 0.085 0.1248 24.86 0.9118 0.087 0.1401
DPCS [38] 25.98 0.9315 0.052 - 25.73 0.9253 0.056 -
GS-ProCams w/o PSF 23.97 0.9222 0.065 0.0725 23.50 0.9100 0.073 0.0630
GS-ProCams (Ours) 26.71 0.9437 0.045 0.0677 26.17 0.9336 0.052 0.0561
50 DeProCams [23] 24.30 0.9143 0.086 0.1451 24.35 0.9074 0.090 0.1454
DPCS [38] 26.26 0.9355 0.050 - 25.98 0.9290 0.054 -
GS-ProCams w/o PSF 24.38 0.9235 0.064 0.0664 23.86 0.9113 0.072 0.0560
GS-ProCams (Ours) 26.65 0.9434 0.046 0.0691 26.09 0.9331 0.053 0.0578
25 DeProCams [23] 23.20 0.9063 0.094 0.1578 23.31 0.8997 0.099 0.1589
DPCS [38] 26.13 0.9339 0.052 - 25.84 0.9275 0.056 -
GS-ProCams w/o PSF 24.14 0.9192 0.066 0.0688 23.65 0.9070 0.074 0.0585
GS-ProCams (Ours) 26.38 0.9404 0.048 0.0702 25.87 0.9302 0.055 0.0596
Table 4: Quantitative comparison of projector compensation on novel viewpoints with novel projections. Results are averaged over 4 scenes. We use bold and underline to highlight the best and second-best results respectively. We recorded the training time and the GPU memory required for training. The training time reported for view-specific methods is the total time consumed across $V$ novel viewpoints (where $V$ is the number of novel viewpoints tested, $V=3$ per scene), reflecting the repeated training process required for each viewpoint. In contrast, the time reported for GS-ProCams is the single average time per scene, as it is a view-agnostic method that performs projector compensation on novel viewpoints without seeing or retraining on them.
[…]
CompenHR [72] 25.81 0.8650 0.145 $2.7\times V$ min 2.5 GB
DeProCams [23] 23.95 0.8364 0.167 $8.8\times V$ min 22.8 GB
DPCS [38] 24.04 0.8427 0.140 $8.3\times V$ min 6.5 GB
GS-ProCams w/o PSF 23.70 0.8482 0.135 8.8 min 3.1 GB
GS-ProCams (Ours) 24.59 0.8652 0.121 10.5 min 3.1 GB
Uncompensated 12.52 0.629 0.315 - -
 View-specific method
@@ 3 Experiments
To demonstrate the advantages of our GS-ProCams, we compared it with several state-of-the-art ProCams simulation and/or projector compensation methods [13, 23, 38, 72]. In our experiments, we employed Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) [83] to evaluate the ProCams simulation and projector compensation quality. Additionally, we statistically examined the training time, memory requirements, and inference speed measured in frames per second (FPS) on a workstation equipped with an Intel Core i9-12900K CPU, an Nvidia RTX 3090 GPU, and 64 GB of RAM.
@@ 3.1 Synthetic dataset
Our GS-ProCams and the state-of-the-art NeRF-based approach [13] (Nepmap for brevity) are both view-agnostic ProCams simulation methods. Both assume a physically static configuration between the projector and the projection surface, enabling viewpoint-agnostic applications. However, key differences exist. Nepmap embeds the projector into a neural reflectance field [6, 64, 84] and relies on a co-located light source under dark-room assumptions to disentangle an object’s geometry and material properties. In contrast, GS-ProCams focuses on modeling the projection surface within the projector FOV, requires no co-located lighting, supports environmental illumination, and employs a joint optimization strategy that eliminates the need for multi-stage training, as summarized in Table˜1.
Despite these differences, we compare with Nepmap on its four synthetic scenes using the pre-trained models. Each scene contains 36 novel viewpoints for evaluation, where each viewpoint corresponds to a unique novel projection. Quantitative and qualitative results are shown in Table˜2 and Fig.˜3, respectively. Clearly, GS-ProCams outperforms Nepmap in ProCams simulation while showing a significant improvement in efficiency, i.e., about 900 times faster in inference speed and only uses 1/10 of GPU memory for training. Unlike Nepmap, GS-ProCams does not disentangle full-object material properties, but focuses on the projection surface within the projector’s FOV. In such scenarios, GS-ProCams interprets co-located lighting as residual color and incorporates projector occlusions into the estimated albedo, as shown in Fig.˜4. Moreover, since synthetic data ignores defocus blur, we use a degraded version of GS-ProCams without PSF modeling (Eq.˜8), denoted as GS-ProCams w/o PSF for comparison.
3.2 Real-world benchmark dataset
We collected a benchmark dataset in real indoor lighting environments. Our projector-camera system consists of a Panasonic Lumix DC-S5 camera and an EPSON CB-965 projector, and an RGB fill light was employed to create various environmental lighting conditions. It is worth noting that, unlike Nepmap [13], our method does not require additional light sources to work, the RGB fill light is only used to change ambient lighting conditions. The projector and camera resolutions are both set to $800\times 800$. Because Nepmap [13] requires a light source to be co-located with the camera and only works for a dark room as described in Section˜3.1, it is incompatible with our real-world benchmark setup. Therefore, we exclude it from real-world comparisons. Instead, we compare with three state-of-the-art view-specific methods [23, 38, 72], using the same projector-camera setup and environmental lighting conditions to ensure a fair comparison. As we will demonstrate, our approach overcomes a fundamental limitation of view-specific baselines, as it generalizes to novel viewpoints while view-specific methods must be retrained for each viewpoint.
We captured 25 training viewpoints per scene for our view-agnostic GS-ProCams. At each viewpoint, we projected one fixed pattern to enhance surface texture for the SfM process [60], along with one uniform black and four sampling patterns for model training. These sampling patterns are natural textured images, following [23]. The plain black projection helps constrain the global illumination residual term by suppressing direct projector influence during joint optimization, while the sampling patterns enable diverse modeling of direct projector effects across views.
Beyond the training set, we collected additional data to evaluate model generalization across novel projections and viewpoints. Specifically, we projected all 100 training patterns onto 5 of the 25 training viewpoints to train view-specific baseline methods. Moreover, we randomly repositioned the camera to capture 8 novel viewpoints beyond the original training set.
At each of these 13 evaluation viewpoints (5 trained + 8 novel), we projected 25 additional novel patterns that were never used during training. Ground-truth geometry was acquired using structured light (SL) [44], followed by manual cleaning of the point clouds. The resulting geometry is in a relative scale since SfM [60] performs self-calibration without physical scale references. Projector FOV masks were obtained following [23]. This dataset establishes a comprehensive benchmark for future research on view-agnostic ProCams simulation under real-world indoor conditions. See supplementary material for more details.
Table 5: ProCams simulation results by our GS-ProCams under different numbers of training viewpoints. We keep the number of textured patterns used for training fixed at 100, while reducing the number of training viewpoints. Note that each viewpoint includes one additional pure black pattern, which also decreases correspondingly. Novel projection synthesis is performed on 8 novel viewpoints. GS-ProCams is capable of maintaining high-quality results even with as few as 4 training viewpoints.
@@ 3.2 Real-world benchmark dataset
    25 26.23 0.9341 0.051 0.0556
    20 26.24 0.9341 0.051 0.0583
    10 26.34 0.9349 0.050 0.0626
    5 26.15 0.9327 0.051 0.0571
    4 26.02 0.9320 0.052 0.0548
    2 24.36 0.9175 0.066 0.0600
Figure 7: Visual comparisons of projector compensation for novel projections. Each row shows a different surface. The first two columns show the projection surfaces and the uncompensated appearance, respectively. The remaining columns display the target appearance and camera-captured compensated results using different methods. Unlike the three baselines that have seen and been retrained on each viewpoint, our GS-ProCams effectively compensates both geometric and photometric distortions, even if it has not seen or been trained on the four novel viewpoints. See supplementary material for more results.
Figure 8: Visual comparison w/ and w/o projector PSF estimation. GS-ProCams w/o PSF produces unrealistically sharp projections inconsistent with the ground truth (first row), and fails on complex geometry with large DoF (second row). In contrast, GS-ProCams yields results that better match real-world observations, enabling more robust ProCams simulation.
ProCams simulation. We use two state-of-the-art methods, DeProCams [23] and DPCS [38] (view-specific), as baseline methods to evaluate our approach (view-agnostic) on the real-world benchmark dataset. We conduct quantitative evaluations by training all models using projector FOV masks for camera viewpoints and computing metrics within the valid SL region, as presented in Table˜3. Our method outperforms both baselines in synthesizing novel projections, not only at training viewpoints but also under previously unseen novel viewpoints, due to our GS-based formulation.
We also illustrate qualitative findings with a scene example in Fig.˜5. We observe that DeProCams may produce blurry results when the projector and viewpoint diverge significantly, potentially due to its reliance on CNNs for shading while simultaneously estimating view-specific per-pixel depth. This poses challenges for high-resolution projections. DPCS, on the other hand, exhibits noise artifacts in some results, likely due to the nature of path tracing [38]. Moreover, DPCS relies on a structured light pre-reconstructed surface mesh from a single viewpoint, while GS-ProCams jointly optimizes geometry across multiple viewpoints. While GS-ProCams performs robustly across different viewpoints, it may introduce inconsistencies in regions outside the projection surface, such as shadowed areas and the background, and minor aliasing artifacts along sharp boundaries. However, GS-ProCams effectively solves the view-agnostic projection mapping problem through physically-based rendering and joint optimization, producing interpretable estimates of the projection surface within the projector FOV from multi-view training images, as shown in Fig.˜6.
Projector compensation. We evaluate the practical effectiveness of our framework on projector compensation, with a particular focus on its ability to generalize to unseen novel viewpoints. We compare our method against three state-of-the-art view-specific methods: CompenHR [72], DeProCams [23], and DPCS [38], all of which require retraining for each specific viewpoint. Since a single projector cannot compensate for regions occluded from the projector direct light, we further evaluate performance on four non-planar, textured surfaces following [72]. We quantitatively evaluate performance by measuring the difference between the camera-captured compensated results and the desired images, with results in Table˜4 and the visual comparisons in Fig.˜7.
The training time reported in Table˜4 for view-specific methods is the total aggregated time consumed across $V$ novel viewpoints ($V=3$ per scene in our setup), reflecting the repeated, per-viewpoint retraining requirement of these approaches. In contrast, our view-agnostic GS-ProCams reports a single average time per scene, highlighting its immediate readiness for deployment upon initial training. The results demonstrate our method’s ability to effectively compensate for both the surface’s geometric distortions and photometric interferences at these novel viewpoints without retraining, all while maintaining substantial advantages in overall training time and GPU memory usage, making it a highly efficient and practical solution for real-world applications.